# Flat-band FFLO State from Quantum Geometric Discrepancy

### v2

@@ IV. FLAT-BAND HELICAL PAIRING STATE AND SUPERCONDUCTING DIODE EFFECT
Here the term with parameter $\gamma$ explicitly breaks the inversion symmetry in the Bloch wave function such that $\alpha_{-\bm{k},\sigma}\neq\alpha_{\bm{k},\sigma}$, while still preserving the time-reversal symmetry as $\alpha_{\bm{k},\uparrow}=\alpha_{-\bm{k},\downarrow}$ when $\eta=0$. However, when $\eta\neq 0$, this term entirely lifts ±$\pm\bm{Q}$ degeneracy in the free energy [see Fig. S4(d)], resulting in FF-type pairing state. Here the Cooper pair momentum is along the $\Gamma-M$ line as $\bm{q}=q(\hat{x},\hat{y})$.
The mean-field results are summarized in Fig. S4. In this case, the finite momentum pairing state is favored for any nonzero $\eta$. Interestingly, $\bm{Q}(\eta)$ shows a sharp change near a critical $\eta_{c}$, manifested as the sign change [Fig. S4 (b)], which suggests a first-order phase transition. In Fig. S4 (a), we color the $\bm{Q}>0$ and $\bm{Q}<0$ regimes in blue and yellow, respectively. This is an analog to the weak and strong helical superconducting states Daido et al. (2022).
@@
We provide some details in evaluating the supercurrent. The supercurrent can be evaluated by $J(q)=2[\partial F(q)]/\partial q$, and the minimized free energy can be calculated from Eq. (S22). To give a clearer picture of the phase transition, in Fig. S4(c) and (d), we show the free energy $F(q)$ (red curves) as well as the supercurrent $J(q)$ (blue curves) for the two points (green and red dots) in Fig. S4(a). In the blue region [Fig. S4 (a)], the global minimum of the free energy $F(q)$ is at $Q>0$ [Fig. S4 (c)], but $|\bm{Q}|$ is small, associated with the small diode efficiency. When $\eta>\eta_{c}$ [Fig. S4 (d)], the pairing state with $Q<0$ becomes the new global minimum. In this case, an enhanced diode efficiency can be observed in Fig. S4(b).

### v3

@@ X Acknowledgements
[…]
Figure S5: (a) The superconducting phase diagram of $\zeta$-lattice model with an inversion breaking term $\gamma$. The $\bm{Q}>0$ (weak helical superconducting state) and $\bm{Q}<0$ (strong helical superconducting state) regimes are colored in blue and yellow, respectively. Parameters used here: $U=0.4t$, $\zeta_{\uparrow}=3$, $\gamma=3$. (b) The red curve shows the finite-momentum $\bm{Q}=Q(1,1)$ versus QGD ($\eta$). Meanwhile, the blue curve is the corresponding diode efficiency $\Theta$. When increasing $\eta$ to across the phase boundary $\eta=0.26$, the diode efficiency $\Theta$ is enhanced dramatically, and the sign of $\bm{Q}$ is reversed. (c) and (d) The free energy (red curve) and supercurrent (blue curve) of 2 highlighted points in (a) with the temperature $T=0.1T_{0}$ but different $\eta$ $\eta=0.24$ in (c) and $0.29$ in (d)]. Both the free energy and the supercurrent are normalized to [-1, 1]. The green point is in the $\bm{Q}>0$ phase, and the red point is in the $\bm{Q}<0$ phase.
The mean-field numerical results are summarized in Fig. S5. In this case, the finite momentum pairing state is favored for any nonzero $\eta$. Interestingly, $\bm{Q}(\eta)$ shows a sharp change near a critical $\eta_{c}$, manifested as the sign change [see Fig. S5 (b)], which suggests a first-order phase transition. In Fig. S5(a), we color the $\bm{Q}>0$ and $\bm{Q}<0$ regions in blue and yellow, respectively. This is an analog to the weak and strong helical superconducting states Daido et al. (2022).
@@
The supercurrent can be evaluated by $J(q)=2[\partial F(q)]/\partial q$, and the minimized free energy can be calculated from Eq. (S29). To give a clearer picture of the phase transition, in Fig. S5(c) and (d), we show the free energy $F(q)$ (red curves) as well as the supercurrent $J(q)$ (blue curves) for the two points (green and red dots) in Fig. S5(a). In the blue region [Fig. S5 (a)], the global minimum of the free energy $F(q)$ is at $Q>0$ [Fig. S5 (c)], but $|\bm{Q}|$ is small, associated with the small diode efficiency. When $\eta>\eta_{c}$ [Fig. S5 (d)], the pairing state with $Q<0$ becomes the global minimum. In this case, an enhanced diode efficiency can be observed in Fig. S5(b).
III.3 C. Inter-flavor FFLO phase within pseudo-Landau levels
In this section, we explore the formation of the FFLO phase in pseudo-Landau level systems. As schematically illustrated in Fig. S6, we consider a strained spinless Dirac material with a series of pseudo-Landau levels under a uniform pseudomagnetic field. The $K$ and $K^{\prime}$ valleys are time-reversal partners, which naturally leads to the inter-flavor Cooper pairings in the presence of attractive interactions. Under this scenario, we consider Cooper pairs form within Landau level $n_{1}$ from $K$ valley and $n_{2}$ from $K^{\prime}$ valley. Here, an imbalance in the populations of the two flavors (such as valley polarization), i.e., $\mu_{K}\neq\mu_{K^{\prime}}$ may lead to $n_{1}\neq n_{2}$, where QGD comes into being as the quantum metrics of the different pseudo-Landau levels are distinct, and the FFLO phase can manifest. This scenario is also proposed in Skyrmion crystals Dong et al. (2024).
To evaluate the critical temperature $T_{cq}$, we first use the form factor of Landau levels on the torus geometry to obtain the anomalous quantum distance Liu et al. (2024):
$\langle u_{n_{2}}^{K}(\bm{k}+\bm{q})|u_{n_{1}}^{*K^{\prime}}(-\bm{k})\rangle=\langle u_{n_{2}}^{K}(\bm{k}+\bm{q})|u_{n_{1}}^{K}(\bm{k})\rangle=e^{\frac{i}{2}(\bm{k}+\bm{q})\times\bm{k}}\sqrt{\frac{n_{1}!}{n_{2}!}}L_{n_{1}}^{n_{2}-n_{1}}(qq^{*})(iq)^{n_{2}-n_{1}}e^{-\frac{1}{2}qq^{*}}.$ (S32)
Here $q=(q_{x}+q_{y})/\sqrt{2}$ and $L_{n}^{m}(x)$ is the generalized Laguerre polynomial. And we have used $|u_{n_{1}}^{*K^{\prime}}(-\bm{k})\rangle=|u_{n_{1}}^{K}(\bm{k})\rangle$. Then we use the formula in main text $T_{c\bm{q}}/T_{0}=1-\bar{d}^{2}_{\mathcal{A}}(\bm{q})$ to calculate $T_{c\bm{q}}$ as summarized in Fig. S6.
Firstly, we notice that the zero-momentum pairing is favored when $n_{1}=n_{2}=n$ as shown in Fig. S6(b). Notice that for Landau levels the quantum metric satisfies $\int tr[g_{n}(\bm{k})]=2n+1$. Thus $T_{cq}/T_{0}=1-\frac{2n+1}{2}q^{2}$ for small $q$. In Fig. S6(c), when $n_{1}\neq n_{2}$, it can be seen that a global maximum occurs at $q=Q$ corresponding to the FFLO phase.
We can make an rough estimation of the finite momentum $Q$ by evaluating $\partial[L_{n_{1}}^{n_{2}-n_{1}}(q^{2}/2)(q/\sqrt{2})^{n_{2}-n_{1}}]/\partial_{q}=0$ at $q=Q$, leading to
$Q^{2}L_{n_{1}-1}^{n_{2}-n_{1}+1}(Q^{2}/2)+(n_{2}-n_{1})L_{n_{1}}^{n_{2}-n_{1}}(Q^{2}/2)=0.$ (S33)
Expanding it to the lowest order of $Q$ and at large $n_{1}$, $n_{2}$, we can make an approximate result as
$Q\approx\sqrt{\frac{2(n_{2}-n_{1})\binom{n_{2}}{n_{1}}}{(n_{2}-n_{1}+2)\binom{n_{2}}{n_{1}-1}}}=\sqrt{\frac{2(n_{2}-n_{1})(n_{2}-n_{1}+1)}{n_{1}(n_{2}-n_{1}+2)}}.$ (S34)
For $n_{2}$ close to $n_{1}$, $Q\sim\sqrt{n_{2}}-\sqrt{n_{1}}$ as demonstrate in Fig. S6(d). We can also observe a direct relationship between $Q$ and the QGD, as the quantum metrics of the different Landau levels are distinct.
Figure S6: (a) The inter-flavor (valley) pairing involves the electrons on Landau level $n_{1}$ from $K$ valley and $n_{2}$ from $K^{\prime}$ valley. (b) The $T_{cq}$ for the BCS-type pairing ($Q=0$) at $n_{1}=n_{2}=n$. (c) Calculated $T_{cq}/T_{0}$ for the FFLO pairng at $n_{1}\neq n_{2}$. (d) The stablized finite momentum $Q$ scales as $\sqrt{n_{2}}-\sqrt{n_{1}}$ at $n_{1}=10$.